# Yoga for Fourier–Mukai partnership

### v1

@@ 3. Preservation of adjointness
$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})$
$\to\mathbf{L}(t^{\prime})^{\ast}(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})).$
As Lemma˜3.2 and affineness is stable under base change, it follows that each vertical morphism in Equation˜3.1 are affine and locally of finite tor-dimension. Also, $(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [Sta25, Tag 0AU1]). Hence, we have a string of isomorphisms
@@
This completes the proof. ∎
@@ 4. Fully faithfulness & equivalences
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then there is a natural isomorphism $\alpha^{K}\colon(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\to\Phi_{K}\circ(t_{1})_{\ast}$.
@@
$\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E)$ $=\mathbf{R}(t_{2})_{\ast}\mathbf{R}(g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
@@
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)\bigg)$
@@
Let $\beta\colon\mathbf{L}t^{\ast}_{1}E\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ be a morphism in $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ for any $A\in D_{\operatorname{qc}}(Y_{2}\times_{S}T)$. Using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$, adjunction tells us there exists a unique morphism $\beta^{\prime}$ making the following diagram commute,
@@
where $\xi^{t_{1}}$ is the counit of derived pullback and pushforward along $t_{1}$. Now, $\beta^{\prime}$ is of the form $\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$ where $\beta^{\prime\prime}\colon E\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$. Furthermore, by the universal property of the counit, the morphism $\beta^{\prime\prime}$ is the unique morphism such that $\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$. Similarly, using that $\Phi_{K}$ is left adjoint to $\Phi_{K^{\prime}}$ and the natural isomorphism $\alpha$ of Lemma˜4.1, there is a $\beta^{\prime\prime\prime}$ giving a diagram
@@
which commutes. In other words,
@@
It remains to show that this factorization is unique. So, suppose there is a morphism $h\colon\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ such that $\beta=h\circ\mathbf{L}t_{1}^{\ast}(\eta_{E})$. Once again, using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$, we have a morphism
@@
However, by the previous work, we know that
@@
Let $F\colon\mathcal{S}_{1}\leftrightarrows\mathcal{S}_{2}\colon G$ be an adjoint pair of exact functors between compactly generated triangulated categories. Assume $G$ preserves small coproducts. Denote by $\eta$ for the unit of the adjunction. Then the collection $\mathcal{T}$ in $\mathcal{S}_{1}$ consisting of objects $E$ for which $\eta_{E}$ is an isomorphism is a triangulated subcategory of $\mathcal{S}_{1}$ which is closed under small coproducts. Moreover, $F$ is fully faithful if, and only if, $\eta_{E}$ is an isomorphism for every $E\in\mathcal{S}_{1}^{c}$.
@@
It suffices to show $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. Note that $\eta_{\oplus_{i\in I}E_{i}}$ is the unique morphism arising from the universal property of $\oplus_{i\in I}E_{i}$ being a coproduct of the $E_{i}$. Hence, if $GF(\oplus_{i}E_{i})$ with morphisms $GF(\pi_{i})\circ\eta_{E_{i}}$ satisfies the necessary universal property, then $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. Consider an $A\in\mathcal{T}$ and morphisms $f_{i}\colon E_{i}\to A$. This gives us morphisms $f_{i}\colon\eta^{-1}_{E_{i}}\colon GF(E_{i})\to A$, and hence, a unique morphism $f\colon\oplus_{i\in I}GF(E_{i})\to A$ such that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\pi^{\prime}_{i}$ for all $i\in I$. It can be checked that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\alpha^{-1}\circ GF(\pi_{i})$, and by uniqueness of $f$, we can see that $f\circ\alpha^{-1}$ is the only morphism from $GF(\oplus_{i\in I}E_{i})$ which can satisfy this property. Consequently, we have $GF(\oplus_{i\in I}E_{i})$ satisfying the necessary universal property to show that $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. So, $\mathcal{T}$ is closed under small coproducts.
@@
Consider ˜3.1 where $t=1_{S}$ and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the right adjoint for $\Phi_{K}$ obtained in Proposition˜3.8. Set $\eta$ to be the unit of the adjunction $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$. Then the following are equivalent:
@@
Lastly, we check the case of $\Phi_{K}$ being an equivalence. From the first claim, it suffices to show $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective. If $\Phi_{K}$ is an equivalence and is left adjoint to $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$, then we can arrange this data so that $\Phi_{K^{\prime}}$ is left adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. By Proposition˜3.8, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Moreover, as $\Phi_{K^{\prime}}$ is an equivalence on $D_{\operatorname{qc}}$, it must also restrict to an equivalence on $\operatorname{Perf}$. Hence, $K^{\prime}$ must be relatively perfect over $Y_{2}$. Now, we may argue as above with the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ to show the unit of this adjunction is an isomorphism. However, this implies $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective, which completes the proof. ∎
@@
Next, we check the case of $\Phi_{K}$ being an equivalence on $D^{b}_{\operatorname{coh}}$. By Remark˜2.7, $\Phi_{K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ implies it restricts to an equivalence on $D_{\operatorname{qc}}$. From Theorem˜4.6, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an equivalence on $D_{\operatorname{qc}}$. However, once again, Remark˜2.7 tells us $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ as desired. ∎
5. Affine variations
This section is concerned with analogs of Sections˜3 and 4 where $t$ in ˜3.1 need not be locally of finite tor-dimension. Instead, we consider the special case that $t$ is an affine morphism to the base with a mild restriction to one of the $Y_{i}\to S$.
Lemma 5.1.
Consider ˜3.1 where $t$ is an affine morphism and $f_{1}$ is locally projective. Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$. In particular, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
Proof.
That the last claim holds follows from [Riz17, Definition 2.11 & Lemma 2.12]. So, we check the first claim. To start, as $K$ is relatively perfect over $Y_{2}$, Lemma˜3.3 tells us that $\mathbf{L}(t^{\prime})^{\ast}K$ is relatively perfect over $Y_{2}\times_{S}T$. Hence, via Lemma˜3.11, we know that $\Phi_{\widetilde{K^{\prime}}}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ where
$\widetilde{K^{\prime}}=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}).$
So, it is enough to show that
$\mathbf{L}(t^{\prime})^{\ast}K^{\prime}=\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}).$
Observe, by [Sta25, Tag 08I3] and Lemma˜3.7, there exists a natural morphism
$\mathbf{L}(t^{\prime})^{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\to\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
$\xrightarrow{\cong}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}\times_{S}T}).$
However, $t$, and hence $t^{\prime}$, is an affine morphism, which ensures their derived pushforwards reflect the zero objects (see e.g. [Sta25, Tag 0BQT]). Hence, it suffices to check that the induced map
$\mathbf{R}t^{\prime}_{\ast}\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\to\mathbf{R}t^{\prime}_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
is an isomorphism after applying $\mathbf{R}t^{\prime}_{\ast}$. Now, we have a string of isomorphisms,
$\mathbf{R}t^{\prime}_{\ast}$ $\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,\mathbf{R}t^{\prime}_{\ast}(\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}))$ ([GW23, Theorem 21.128])
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,\mathbf{R}t^{\prime}_{\ast}\mathcal{O}_{Y_{1}\times_{S}Y_{2}\times_{S}T}\otimes^{\mathbf{L}}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$ (Projection formula)
$\cong\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\otimes\mathbf{L}(f_{1}^{\prime})^{\ast}\mathbf{R}(t_{2})_{\ast}\mathcal{O}_{Y_{2}\times T})$ $\textrm{(Flat base change)}.$
Then we have a commutative square
$\mathbf{R}t^{\prime}_{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,\mathbf{L}(t^{\prime})^{\ast}(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$$\mathbf{R}t^{\prime}_{\ast}\mathbf{L}(t^{\prime})^{\ast}\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})\otimes\mathbf{L}(f_{1}^{\prime})^{\ast}\mathbf{R}(t_{2})_{\ast}\mathcal{O}_{Y_{2}\times T}$$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\otimes\mathbf{L}(f_{1}^{\prime})^{\ast}\mathbf{R}(t_{2})_{\ast}\mathcal{O}_{Y_{2}\times T})$$\scriptstyle{\cong}\scriptstyle{\cong}$
where the bottom vertical morphism is an isomorphism (see [Riz17, Lemma 2.15]). This gives us the required isomorphism and completes the proof. ∎
Lemma 5.2.
Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated. Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$. Then the right adjoint of $\Phi_{K}$ restricts to an exact functor $D^{b}_{\operatorname{coh}}(Y_{2})\to D^{b}_{\operatorname{coh}}(Y_{1})$. A similar statement holds for the right adjoint of $\Phi_{\mathbf{L}(t^{\prime})K}$.
Proof.
The first claim is exactly Corollary˜3.9. By Lemma˜3.3, our hypothesis implies $\mathbf{L}(t^{\prime})K$ is relatively perfect over each $Y_{i}\times_{S}T$. So, once again, the second claim follows from Corollary˜3.9. ∎
Proposition 5.3.
Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ the kernel of the integral transform associated to the right adjoint of $\Phi_{K}\colon D_{\operatorname{qc}}(Y_{1})\to D_{\operatorname{qc}}(Y_{2})$ obtained in Proposition˜3.8. If $\Phi_{K}$ is fully faithful on $D_{\operatorname{qc}}$, then $\mathbf{L}t^{\ast}_{1}$ preserves the unit $\eta$ of the adjoint pair $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$. That is, $\mathbf{L}t_{1}^{\ast}\eta_{E}$ is the unit of $\mathbf{L}t_{1}^{\ast}E$ for the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$.
Proof.
The claims follow essentially verbatim to the proof of Proposition˜4.2 if coupled with Lemmas˜5.1 and 5.2. ∎
Theorem 5.4.
Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$. If $\Phi_{K}$ is fully faithful (resp. an equivalence) on $D_{\operatorname{qc}}$, then so is $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
We can argue as in Theorem˜4.6 by replacing Proposition˜3.8 with Lemma˜5.1 and Proposition˜4.2 with Proposition˜5.3. ∎
Corollary 5.5.
Consider ˜3.1 where $t$ is affine, $f_{1}$ is locally projective, and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
Proof.
This is argued similar to Corollary˜4.7 if coupled with Theorem˜5.4. ∎

### v2

@@ 3. Preservation of adjointness
$\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(\mathbf{L}(t^{\prime})^{\ast}K,(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}\times_{S}T})\to\mathbf{L}(t^{\prime})^{\ast}(\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})).$
As Lemma˜3.2 and affineness is stable under base change, it follows that each vertical morphism in Diagram˜3.1 are affine and locally of finite tor-dimension. Also, $(g^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [Sta25, Tag 0AU1]). Hence, we have a string of isomorphisms
@@
in $D_{\operatorname{coh}}^{+}$. This completes the proof. ∎
@@ 4. Fully faithfulness & equivalences
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then there is a natural isomorphism $\alpha^{K}\colon\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\to\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$.
@@
$\mathbf{R}(t_{2})_{\ast}$ $\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E)$
$=\mathbf{R}(t_{2})_{\ast}\mathbf{R}(g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
@@
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)\bigg)$ ([GW23, Proposition 22.84])
@@
Let $\beta\colon\mathbf{L}t^{\ast}_{1}E\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ be a morphism in $D_{\operatorname{qc}}(Y_{1}\times_{S}T)$ for any $A\in D_{\operatorname{qc}}(Y_{2}\times_{S}T)$. Using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$ on $D_{\operatorname{qc}}$, adjunction tells us there exists a unique morphism $\beta^{\prime}$ of the form $\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$ where $\beta^{\prime\prime}\colon E\to\mathbf{R}(t_{1})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$, making the following diagram commute,
@@
where $\xi^{t_{1}}$ is the counit of derived pullback and pushforward along $t_{1}$. Furthermore, by the universal property of the counit, the morphism $\beta^{\prime\prime}$ is the unique morphism such that $\beta=\xi^{t_{1}}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)}\circ\mathbf{L}t^{\ast}_{1}(\beta^{\prime\prime})$. Similarly, using that $\Phi_{K}$ is left adjoint to $\Phi_{K^{\prime}}$ and the natural isomorphism $\alpha$ of Lemma˜4.1, there is a $\beta^{\prime\prime\prime}$ giving a diagram
@@
which commutes. Pulling back this diagram via $\mathbf{L}t_{1}^{\ast}$, we obtain
@@
It remains to show that the above factorization for $\beta^{\prime}$ is unique. So, suppose there is a morphism $h\colon\mathbf{L}t_{1}^{\ast}\Phi_{K^{\prime}}\Phi_{K}(E)\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}(A)$ such that $\beta=h\circ\mathbf{L}t_{1}^{\ast}(\eta_{E})$. Once again, using that $\mathbf{L}t_{1}^{\ast}$ is left adjoint to $\mathbf{R}(t_{1})_{\ast}$, we have a morphism
@@
However, by the work above, we know that
@@
Let $F\colon\mathcal{S}_{1}\leftrightarrows\mathcal{S}_{2}\colon G$ be an adjoint pair of exact functors between compactly generated triangulated categories. Assume $G$ preserves small coproducts. Denote by $\eta\colon 1\to GF$ for the unit of the adjunction. Then the collection $\mathcal{T}$ in $\mathcal{S}_{1}$ consisting of objects $E$ for which $\eta_{E}$ is an isomorphism is a triangulated subcategory of $\mathcal{S}_{1}$ which is closed under small coproducts. Moreover, $F$ is fully faithful if, and only if, $\eta_{E}$ is an isomorphism for every $E\in\mathcal{S}_{1}^{c}$.
@@
It suffices to show $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. Note that $\eta_{\oplus_{i\in I}E_{i}}$ is the unique morphism arising from the universal property of $\oplus_{i\in I}E_{i}$ being a coproduct of the $E_{i}$. Hence, if $GF(\oplus_{i}E_{i})$ with morphisms $GF(\pi_{i})\circ\eta_{E_{i}}$ satisfies the necessary universal property, then $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. Consider an $A\in\mathcal{T}$ and morphisms $f_{i}\colon E_{i}\to A$. This gives us morphisms $f_{i}\circ\eta^{-1}_{E_{i}}\colon GF(E_{i})\to A$, and hence, a unique morphism $f\colon\oplus_{i\in I}GF(E_{i})\to A$ such that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\pi^{\prime}_{i}$ for all $i\in I$. It can be checked that $f_{i}\circ\eta^{-1}_{E_{i}}=f\circ\alpha^{-1}\circ GF(\pi_{i})$, and by uniqueness of $f$, we can see that $f\circ\alpha^{-1}$ is the only morphism from $GF(\oplus_{i\in I}E_{i})$ which can satisfy this property. Consequently, we have $GF(\oplus_{i\in I}E_{i})$ satisfying the necessary universal property to show that $\eta_{\oplus_{i\in I}E_{i}}$ is an isomorphism. So, $\mathcal{T}$ is closed under small coproducts.
@@
Consider ˜3.1 where $t=1_{S}$ and $S$ is semiseparated. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the right adjoint for $\Phi_{K}$ obtained in Proposition˜3.8. Set $\eta\colon 1\to\Phi_{K}\Phi_{K^{\prime}}$ to be the unit of the adjunction $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$. Then the following are equivalent:
@@
Lastly, we check the case of $\Phi_{K}$ being an equivalence. From the first claim, it suffices to show $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective. If $\Phi_{K}$ is an equivalence and is left adjoint to $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$, then one can show that $\Phi_{K^{\prime}}$ is also left adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. By Proposition˜3.8, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$. Moreover, as $\Phi_{K^{\prime}}$ is an equivalence on $D_{\operatorname{qc}}$, it must also restrict to an equivalence on $\operatorname{Perf}$. Hence, $K^{\prime}$ must be relatively perfect over $Y_{1}$. Now, by the above $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is fully faithful and by Proposition 4.4 the unit of the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an isomorphism. However, this implies $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is essentially surjective, which completes the proof. ∎
@@
Next, we check the case of $\Phi_{K}$ being an equivalence on $D^{b}_{\operatorname{coh}}$. By Remark˜2.7, $\Phi_{K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ implies $\Phi_{K}$ is an equivalence on $D_{\operatorname{qc}}$. From Theorem˜4.6, $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ is an equivalence on $D_{\operatorname{qc}}$. However, once again, Remark˜2.7 tells us $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an equivalence on $D^{b}_{\operatorname{coh}}$ as desired. ∎